# The MUSE eXtremely Deep Field: Detections of circumgalactic Siāii*Ā emission at ā ā thanks: Based on observations made with ESO telescope at the La Silla Paranal Observatory under the large program 1101.A-0127

### v1

@@ 2.2 Sample selection
We construct a sample of galaxies at ā, for which the [Cāiii]1907, Cāiii]1909 doublet nebular lines and at least one of the Siāii*1265, 1309, 1533 lines are redshifted into the MUSE wavelength range. The available redshift ranges for the three Siāii*Ā lines are shown in the top panel in Figure 2. We require sources to have a secure spectroscopic redshift with and a Cāiii] signal-to-noise ratio (SNR) of . We visually inspect Cāiii] in spectra and images extracted from MUSE cubes to ensure these criteria. We use systemic redshifts measured by the Cāiii] line rather than reference redshifts (REFZs) in the DR2 catalog, which are given by Ly at or absorption lines at . Siāii*Ā lines are expected to be located at the systemic redshifts (e.g., France etĀ al. 2010; Dessauges-Zavadsky etĀ al. 2010; Jaskot & Oey 2014; Wang etĀ al. 2020), though redshifted Siāii*Ā lines are also reported, which could be due to the contamination of the absorption at the blue edge for low-resolution spectra or stacking of spectra for different sources (e.g., Shapley etĀ al. 2003; Erb etĀ al. 2010; Berry etĀ al. 2012). As Siāii*Ā lines are extremely faint, secure measurements of systemic redshifts enable us to not rely on the presence of Siāii*Ā in 1D spectra on a galaxy scale and to directly search for spatially extended Siāii*Ā in narrowband images extracted from MUSE cubes. We limit our sample to galaxies with a secure and isolated HST counterpart (based on Rafelski etĀ al. 2015, MAG_SRCRAF) with a high HST matching confidence level of MCONF in the DR2 catalog. In total, we have 39 galaxies, whose absolute UV magnitudes () and systemic redshifts () are shown in the bottom panel of Figure 2. The is calculated at rest-frame 1600 Ā Ć Ā by fitting 2 or 3 HST bands with a power-low model (with parameters being an amplitude and a UV slope ).
@@ 2.3 Stacked sample
The subsample used for the stacking analysis is restricted to bright UV-continuum galaxies using an apparent magnitude cut of , in order to enhance the SNR of the stacked images (see Section 3.3 and Figure 2). The origin of Siāii*Ā is predicted to be continuum pumping as mentioned in Section 1, and a bright UV continuum is necessary to have a bright Siāii*Ā emission line. Indeed, Siāii*Ā emission lines are stronger in 1D stacked spectra for subsamples of MUSE LAEs with brighter UV continuum (such as those with higher stellar masses and with brighter UV magnitudes) than the subsample counterparts in Feltre etĀ al. (2020). The number of sources for the stacking subsample is 14. We would like to note that applying fainter magnitude cuts indeed reduces the SNRs of the Siāii*Ā lines.
@@ 2.4 Masks
In order to exclude pixels that might be affected by bright neighboring objects in Siāii*Ā narrowband images extracted from the minicubes, we create neighboring object masks. First, we define target pixels. We use an HST segmentation map and mask pixels not corresponding to a main target on a 5ā5ā HST/F775W cutout for each object. It is convolved with the MUSE moffat PSF at the redshifted Siāii*Ā wavelength for the main target and rebinned to match the MUSE pixel scale. Then, we apply a threshold value of 0.1 to the peak-normalized convolved cutout to delimit the spaxels of the target. Lower threshold values such as 0.05 cause contamination of neighboring objects inside the target spaxels. Using a higher threshold value of 0.2 does not change the main results in this paper but enhances the contaminated pixels outside the following neighboring object masks. Therefore, we adopt the value 0.1.
Second, we use a 5ā5ā HST/F775W cutout whose sky and the main target are masked with the HST segmentation map for each source. This cutout is convolved with the MUSE PSF and rebinned to match the MUSE pixel scale. Then, we define pixels brighter than a certain threshold on the cutout, except for the target pixels, as masked pixels. The applied threshold value is half of the value above in the absolute sense (without normalization).
@@ 3.1 Extracted images and surface brightness profiles
For each source, we create a Siāii*Ā emission narrowband (NB) image from the continuum-subtracted minicube. We sum fluxes in a window of 200 km s-1 around the Siāii*Ā wavelength, excluding NaN (masked) spaxels, wavelength slices affected by OH skylines, and the AO gap in the cube. The variance image of the NB is created from that of the cube with error propagation. A UV continuum broadband (BB) is created from the median filtered original minicube with a window of 300 wavelength slices (375Ā Ć ) around the Siāii*Ā wavelength (masking below Ā kmĀ s-1 from the Ly wavelength) to obtain a good SNR. We also exclude NaN in the cube and the AO gap. If the wavelength window is not fully covered, we use as many spectral slices as possible. The variance image of the BB is assumed to be identical to that of the mean-filtered BB created from the same cube with error propagation. We apply a neighboring object mask (see Section 2.4) and measure SB profiles of these two bands using PHOTUTILS (Bradley etĀ al. 2021). The aperture and annuli are centered at the HST counterpart position with 1-pixel () annuli from to and then 2-pixel annuli at and , followed by a 3.5-pixel annulus at , to enhance the SNR at outer radii. The azimuthally-averaged radial SB profiles are calculated from the effective areas and the measured flux above. We normalize the SB profile of the UV continuum at the center () to that of Siāii*.
@@
The top panels in Figure 3 show the Siāii*Ā emission NB, the UV-continuum BB, and the SB profile for a highlighted object, MID. The details of this object are explained later (Section 3.2).
3.2 Identification of Siāii*Ā halos
@@
To identify extended Siāii*Ā emission, we compare the shape of the azimuthally-averaged radial SB profiles of the UV continuum and the Siāii*Ā emission. We take account of uncertainties in SB profiles of Siāii*Ā emission only (blue shades in Figure 3) as those of UV continuum are negligibly small (gray shades). If more than two adjacent data points of the Siāii*Ā SB profiles deviate by more than from the normalized continuum, we identify it as extended Siāii*. The significance of the presence of the halo with this criterion corresponds to more than 97.5%333The probability for two adjacent data points with , which is caused by chance due to noises (i.e., fake detections), is =2.5% with an assumption that noises follow the normal distribution.. We would like to note that, as we use circular photometry, our test may miss non-circular halos as discussed in Kusakabe etĀ al. (2022) for Ly halos. Our halo search is not complete in that sense but secure.
@@
Among the 39 sources, we detect five Siāii*Ā halos for the Siāii* 1533 line (MID, , , , ; see Figure 3). These are the first detections of individual Siāii*Ā halos. The Siāii*Ā lines in the DR2 reference spectra show that there are no significant velocity shifts for the line peaks of Siāii*1533 (Figure 3; the systemic redshifts are given by the [Cāiii]1907, Cāiii]1909 doublet nebular lines, see Section 2.2).
@@
Among the five Siāii*Ā halo sources, MID=, shown in the top panels of Figure 3, is the best-case object. The Siāii*Ā emission profile shows three data points that deviate by more than from the normalized continuum. This object is well-isolated and without a spatial offset of Siāii*Ā from the UV continuum. The images are not significantly affected by masks, which makes the SB profile test secure. The Siāii*Ā emission line can be clearly confirmed on the spectrum shown in Figure 3. MID and MID= are a pair of galaxies, which could contaminate each otherās 5ā 5ā images despite the neighboring object mask. Moreover, MID= is an X-ray detected AGN (Luo etĀ al. 2017), whose Siāii*Ā halo might have a different origin from other sources. The Siāii*Ā NB of MID= shows a peak offset from the UV center measured with the HST data, with two neighboring LAEs in the DR2 catalog, which spatially overlap with the extended Siāii*Ā emission. The Siāii*Ā SB of MID= is noisy even though it meets our halo criterion. Considering these details on the five sources, we highlight MID= in the following sections.
The redshifts of the individually-detected Siāii*1533 halos range from 2.23 to 3.32 with four of them at . The non detections at could be due to the high noise levels of MUSE data at red wavelengths (severe night sky and low sensitivity; Bacon etĀ al. 2023) and the cosmic dimming effect on the SB profiles (by a factor of 2.7 between and ).
@@
Another interesting result is that we cannot confirm extended emission for any of the other Siāii*Ā lines (1265, 1309). In the optically thin limit, the strengths of fluorescent lines are expected to depend on the strength of the paired absorption lines, which are determined by their oscillator strengths, . In particular, the Siāii resonant absorption line corresponding to the Siāii*1265 line has a ten times larger than that of the Siāii resonant absorption line corresponding to the Siāii*1533. However, in the simulations of Mauerhofer etĀ al. (2021) there is no significant difference in the equivalent widths of Siāii*1265 and Siāii*1533 absorption lines because both lines are saturated (Mauerhofer et al. in prep.). In such an optically thick regime, the equivalent widths for the corresponding fluorescent emission (Siāii*1265 and Siāii*1533) are similar in the simulations. Therefore, non detections of Siāii*1265 as well as Siāii*1309 would not be due to fainter total fluxes for these lines than for the 1533 line. The reason for nondetections of Siāii*1265 and Siāii*1309 halos are not clear. The samples for Siāii*1265, 1309 lines are limited to high redshifts ( and , respectively), and none of them are bright in apparent UV magnitudes, which are directly related to SNRs (see Figure 2). Bright sources would be crucial for the detection of diffuse Siāii*Ā lines if the origin of Siāii*Ā is continuum pumping. However, we cannot distinguish whether it is caused by differences in lines, samples, or noise levels. We note that the Siāii* 1533 lines are also brighter than the other two lines in the stacked MUSE LAE spectra of Feltre etĀ al. (2020) and Kramarenko etĀ al. (2023), which might also be due to the difference in objects stacked for different Siāii*Ā lines.
@@
One of the individually-detected objects, MID= at , has 3 Siāii*Ā lines covered with MUSE. We coadd the NBs and BBs for the 3 Siāii*Ā lines by taking the median (with the same method as used in Section 3.3) and show them in the fifth row of Figure 3. Extended Siāii*Ā is confirmed with a better SNR, which implies the Siāii*1265, 1309 lines are also spatially extended.
@@ 3.3 Stacking analysis
To test the general presence of Siāii*Ā halos, we stack 14 UV-bright sources (see Section 2.3). The numbers of sources with the magnitude cut of available for each of the Siāii*1265, 1309, 1533 lines are five, four, and 13, respectively, which are limited by the MUSE wavelength coverage and the AO gap (see Figure 2)444The MIDs for the Siāii*1265 subsample are 50, 76, 6666, 6700, and 8048. Those for the Siāii*1309 are 50, 76, 6700, and 8048. Those for the Siāii*1533 are 24, 35, 50, 51, 76, 1141, 1316, 1395, 6664, 6666, 6700, 7600, and 8513.. Therefore, we only stack a subsample for the Siāii*1533 line and take the mean and median of the Siāii*Ā NBs and BBs without weighting, while masking neighboring objects. We also stack 14 sources by combining the NBs and BBs for the three lines. The variance images for the mean stacking, which we use for our halo test, are created based on error propagation. Then, we apply the SB profile test for the mean-stacked images with the same method as used in Sections 3.1 and 3.2, as the uncertainties can be calculated correctly for the mean but not for the median.
@@
The results of this stacking experiment are shown in Figure 4. As shown in the top row, we confirm the existence of a Siāii*1533 halo with five data points that deviate by more than from the normalized continuum profile. The median profile of Siāii*1533 is consistent with the mean profile and deviates from the median continuum as well, in particular at large radii. These tests may suggest that the detected halo in the mean profile is not dominated by a few extreme objects. This is the first secure detection of a stacked Siāii*Ā halo. It may imply that the individually detected sources are not extreme cases, Siāii*Ā halos must be common for UV-bright galaxies at ā.
@@
The bottom row of Figure 4 shows the results of stacking all the Siāii*Ā narrowbands and corresponding UV continuum bands for the 14 sources (three lines; in total 22 images). We find that the mean-stacked Siāii*Ā deviates by more than from the mean-stacked UV continuum for six data points. The median-stacked Siāii*Ā is consistent with the mean-stacked Siāii*, and the deviation from the normalized continuum is similarly seen with the median-stacked profiles.
@@
We note that removing MID, which is an X-ray detected AGN with extended Siāii*Ā (see Section 3.2), does not change the results and that the other 13 sources do not have a counterpart in the deep X-ray catalog (Luo etĀ al. 2017). We also note that stacking without individually-detected Siāii*Ā halos results in nondetections of extended Siāii*. As shown in Figure 2, all the individually-detected sources have a bright (and also a bright ) and are located at relatively low redshifts. It is not surprising that removing such five sources results in nondetections. If the origin of the Siāii*Ā is continuum pumping, as we expect, it then should be more challenging to detect Siāii*Ā halos without these bright sources. We also stack various subsamples including all the sources without individually-detected halos and get nondetections of extended Siāii*Ā emission. Future instruments such as BlueMUSE (Richard etĀ al. 2019) or a larger sample compiling MUSE and KCWI archive data will allow us to investigate the dependence of Siāii*Ā halos on sample properties such as (see Section 4.3). It will enable us to draw a firm conclusion of whether Siāii*Ā halos are indeed common for UV-bright galaxies at ā or not.

### v2

@@ 2.2 Sample selection
We construct a sample of galaxies at –, for which the [C iii]1907, C iii]1909 doublet nebular lines and at least one of the Si ii* 1265, 1309, 1533 lines are redshifted into the MUSE wavelength range. The available redshift ranges for the three Si ii* lines are shown in the top panel in Fig. 2. We require sources to have a secure spectroscopic redshift with and a C iii] signal-to-noise ratio (S/N) of . We visually inspect C iii] in spectra and images extracted from MUSE cubes to ensure these criteria. We use systemic redshifts measured by the C iii] line rather than reference redshifts (REFZs) in the DR2 catalog, which are given by Ly at or absorption lines at . Si ii* lines are expected to be located at the systemic redshifts (e.g., France et al. 2010; Dessauges-Zavadsky et al. 2010; Jaskot & Oey 2014; Wang et al. 2020), though redshifted Si ii* lines are also reported, which could be due to the contamination of the absorption at the blue edge for low-resolution spectra or stacking of spectra for different sources (e.g., Shapley et al. 2003; Erb et al. 2010; Berry et al. 2012). As Si ii* lines are extremely faint, secure measurements of systemic redshifts enable us to not rely on the presence of Si ii* in 1D spectra on a galaxy scale and to directly search for spatially extended Si ii* in narrowband images extracted from MUSE cubes. We limit our sample to galaxies with a secure and isolated HST counterpart (based on Rafelski et al. 2015, MAG_SRCRAF) with a high HST matching confidence level of MCONF in the DR2 catalog. In total, we have 39 galaxies, which are summarized in Table 1. Absolute UV magnitudes () and systemic redshifts () are shown in the middle panel of Fig. 2. The is calculated at rest-frame 1600 Å by fitting 2 or 3 HST bands with a power-low model (with parameters being an amplitude and a UV slope ). Figure 2 also indicates the redshift coverage of the three Si ii* lines (top panel) and the surface brightness limits for each field (bottom panel). The stellar masses and star formation rates (SFRs) of the sample are shown in Fig. 9 in Appendix.
@@ 2.3 Stacked sample
The subsample used for the stacking analysis is restricted to bright UV-continuum galaxies using an apparent magnitude cut of , in order to enhance the S/N of the stacked images (see Section 3.3 and Fig. 2). The origin of Si ii* is predicted to be continuum pumping as mentioned in Section 1, and a bright UV continuum is necessary to have a bright Si ii* emission line. Indeed, Si ii* emission lines are stronger in 1D stacked spectra for subsamples of MUSE LAEs with brighter UV continuum (such as those with higher stellar masses and with brighter UV magnitudes) than the subsample counterparts in Feltre et al. (2020). The number of sources for the stacking subsample is 14. We would like to note that applying fainter magnitude cuts indeed reduces the S/Ns of the Si ii* lines.
@@ 2.4 Masks
In order to exclude pixels that might be affected by bright neighboring objects in Si ii* narrowband images extracted from the minicubes, we create neighboring object masks. First, we define target pixels. We use an HST segmentation map and mask pixels not corresponding to a main target on a 5”5” HST/F775W cutout for each object. It is convolved with the MUSE moffat PSF at the redshifted Si ii* wavelength for the main target and rebinned to match the MUSE pixel scale. Then, we apply a threshold value of 0.1 to the peak-normalized convolved cutout to delimit the spaxels of the target. Lower threshold values such as 0.05 cause contamination of neighboring objects inside the target spaxels. Using a higher threshold value of 0.2 does not change the main results in this paper but enhances the contaminated pixels outside the following neighboring object masks. Therefore, we adopt the value 0.1.
Second, we use a 5”5” HST/F775W cutout whose sky and the main target are masked with the HST segmentation map for each source. This cutout is convolved with the MUSE PSF and rebinned to match the MUSE pixel scale. Then, we define pixels brighter than a certain threshold on the cutout, except for the target pixels, as masked pixels. The applied threshold value is half of the value above in the absolute sense (without normalization).
@@ 3.1 Extracted images and surface brightness profiles
For each source, we create a Si ii* emission narrowband (NB) image from the continuum-subtracted minicube. We sum fluxes in a window of 200 km s-1 around the Si ii* wavelength, excluding NaN (masked) spaxels, wavelength slices affected by OH skylines, and the AO gap in the cube. The variance image of the NB is created from that of the cube with error propagation. A UV continuum broadband (BB) is created from the median filtered original minicube with a window of 300 wavelength slices (375 Å) around the Si ii* wavelength (masking below km s-1 from the Ly wavelength) to obtain a good S/N. We also exclude NaN in the cube and the AO gap. If the wavelength window is not fully covered, we use as many spectral slices as possible. The variance image of the BB is assumed to be identical to that of the mean-filtered BB created from the same cube with error propagation. We apply a neighboring object mask (see Section 2.4) and measure SB profiles of these two bands using PHOTUTILS (Bradley et al. 2021). The aperture and annuli are centered at the HST counterpart position with 1-pixel () annuli from to and then 2-pixel annuli at and , followed by a 3.5-pixel annulus at , to enhance the S/N at outer radii. The azimuthally-averaged radial SB profiles are calculated from the effective areas and the measured flux above. We normalize the SB profile of the UV continuum at the innermost radius (R=) to that of Si ii*.
@@
The top panels in Fig. 3 show the Si ii* emission NB, the UV-continuum BB, and the SB profile for a highlighted object, MID. The details of this object are explained later (Section 3.2).
3.2 Identification of Si ii* halos
@@
To identify extended Si ii* emission, we compare the shape of the azimuthally-averaged radial SB profiles of the UV continuum and the Si ii* emission. We take account of uncertainties in SB profiles of Si ii* emission only (blue shades in Fig. 3) as those of UV continuum are negligibly small (gray shades). If more than two adjacent data points of the Si ii* SB profiles deviate by more than from the normalized continuum, we identify it as extended Si ii*. The significance of the presence of the halo with this criterion corresponds to more than 97.5%444The probability for two adjacent data points with , which is caused by chance due to noises (i.e., fake detections), is =2.5% with an assumption that noises follow the normal distribution.. We would like to note that, as we use circular photometry, our test may miss non-circular halos as discussed in Kusakabe et al. (2022) for Ly halos. Developing and testing non-circular methods requires high-S/N objects for reliable assessments, which our sample is not eligible for. Our halo search is not complete in that sense but secure.
@@
Among the 39 sources, we detect five Si ii* halos for the Si ii* 1533 line (MID, , , , ; see Fig. 3). These are the first detections of individual Si ii* halos. The Si ii* lines in the DR2 reference spectra show that there are no significant velocity shifts for the line peaks of the Si ii* (Fig. 3; the systemic redshifts are given by the [C iii] 1907, C iii] 1909 doublet nebular lines, see Section 2.2). Meanwhile, peaks of the Si ii absorption lines show velocity offsets, which have often been observed in the down-the-barrel analysis and interpreted as gas flows (e.g., Du et al. 2018)555Si ii line is a commonly used tracer of outflows seen along the line of sight. Due to the resonant scattering involved in the 1527 transition (see Fig. 1), observed Si ii* has information on the other line of sight directions. Therefore, these two lines may exhibit different velocity offsets, even if the origin of Si ii* is continuum pumping influenced by resonant scattering..
@@
Among the five Si ii* halo sources, MID=, shown in the top panels of Fig. 3, is the best-case object. The Si ii* emission profile shows three data points that deviate by more than from the normalized continuum. This object is well-isolated and without a spatial offset of Si ii* from the UV continuum. The images are not significantly affected by masks, which makes the SB profile test secure. The Si ii* emission line can be clearly confirmed on the spectrum shown in Fig. 3. MID and MID= are a pair of galaxies, which could contaminate each other’s 5” 5” images despite the neighboring object mask. Moreover, MID= is an X-ray detected AGN (Luo et al. 2017), whose Si ii* halo might have a different origin from other sources. The Si ii* NB of MID= shows a peak offset from the UV center measured with the HST data, with two neighboring LAEs in the DR2 catalog, which spatially overlap with the extended Si ii* emission. The Si ii* SB of MID= is noisy even though it meets our halo criterion. Considering these details on the five sources, we highlight MID= in the following sections.
As shown in Fig. 2, the five individual halos tend to have bright UV magnitudes, which is qualitatively consistent with the continuum pumping scenario of Si ii* emission. We also discuss possible differences in SFRs and between Si ii* 1533 halo sources and nondetected sources in Fig. 9, and also examine the dependence of detections of halos on 1 SB limits for Si ii* NBs in Fig. 10, although they are beyond the scope of this paper to conclude them. Overall, we cannot conclude them due to the small sample size, but UV magnitudes and redshifts of the sources might be essential.
The redshifts of the individually-detected Si ii* 1533 halos range from 2.23 to 3.32 with four of them at . The nondetections at would be caused by the lack of UV-bright sources (see Figs. 2 and 10) if the origin of Si ii* is continuum pumping. It could also be due to the high noise levels of MUSE data at red wavelengths (severe night sky and low sensitivity; Bacon et al. 2023) and the cosmic dimming effect on the SB profiles (by a factor of 2.7 between and ). In case the circumgalactic Si ii*is intrinsically fainter at than at , it implies a rapid evolution of metal enrichment in the CGM, which is qualitatively consistent with the redshift evolution of the number density of Mg ii absorbers (Matejek & Simcoe 2012; Zhu & Ménard 2013) as well as the prediction from a chemical evolution model with UniverseMachine (Behroozi & Silk 2018; Nishigaki et al. in prep.).
@@
Another interesting result is that we cannot confirm extended emission for any of the other Si ii* lines (1265, 1309). In the optically thin limit, the strengths of fluorescent lines are expected to depend on the strength of the paired absorption lines, which are determined by their oscillator strengths, . In particular, the Si ii resonant absorption line corresponding to the Si ii* 1265 line has a ten times larger than that of the Si ii resonant absorption line corresponding to the Si ii* 1533. However, in the simulations of Mauerhofer et al. (2021) there is no significant difference in the equivalent widths of Si ii* 1265 and Si ii* 1533 absorption lines because both lines are saturated (Mauerhofer et al. in prep.). In such an optically thick regime, the equivalent widths for the corresponding fluorescent emission (Si ii* 1265 and Si ii* 1533) are similar in the simulations. Therefore, non detections of Si ii* 1265 as well as Si ii* 1309 would not be due to fainter total fluxes for these lines than for the 1533 line. The reason for nondetections of Si ii* 1265 and Si ii* 1309 halos are not clear. The samples for Si ii* 1265, 1309 lines are limited to high redshifts ( and , respectively), and none of them are bright in apparent UV magnitudes, which are directly related to S/Ns (see Fig. 2). Bright sources would be crucial for the detection of diffuse Si ii* lines if the origin of Si ii* is continuum pumping. However, we cannot distinguish whether it is caused by differences in lines, samples, or noise levels. We note that the Si ii* 1533 lines are also brighter than the other two lines in the stacked MUSE LAE spectra of Feltre et al. (2020) and Kramarenko et al. (2024), which might also be due to the difference in objects stacked for different Si ii* lines.
@@
One of the individually-detected objects, MID= at , has 3 Si ii* lines covered with MUSE. We coadd the NBs and BBs for the 3 Si ii* lines by taking the median (with the same method as used in Section 3.3) and show them in the fifth row of Fig. 3. Extended Si ii* is confirmed with a better S/N, which implies the Si ii* 1265, 1309 lines are also spatially extended.
@@ 3.3 Stacking analysis
To test the general presence of Si ii* halos, we stack 14 UV-bright sources (see Section 2.3). The numbers of sources with the magnitude cut of available for each of the Si ii* 1265, 1309, 1533 lines are five, four, and 13, respectively, which are limited by the MUSE wavelength coverage and the AO gap (see Fig. 2)666The MIDs for the Si ii* 1265 subsample are 50, 76, 6666, 6700, and 8048. Those for the Si ii* 1309 are 50, 76, 6700, and 8048. Those for the Si ii* 1533 are 24, 35, 50, 51, 76, 1141, 1316, 1395, 6664, 6666, 6700, 7600, and 8513.. Therefore, we only stack a subsample for the Si ii* 1533 line and take the mean and median of the Si ii* NBs and BBs without weighting, while masking neighboring objects. We also stack 14 sources by combining the NBs and BBs for the three lines. The variance images for the mean stacking, which we use for our halo test, are created based on error propagation. Then, we apply the SB profile test for the mean-stacked images with the same method as used in Sections 3.1 and 3.2, as the uncertainties can be calculated correctly for the mean but not for the median.
@@
The results of this stacking experiment are shown in Fig. 4. As shown in the top row, we confirm the existence of a Si ii* 1533 halo with five data points that deviate by more than from the normalized continuum profile. The median profile of Si ii* 1533 is consistent with the mean profile and deviates from the median continuum as well, in particular at large radii. These tests may suggest that the detected halo in the mean profile is not dominated by a few extreme objects. This is the first secure detection of a stacked Si ii* halo. It may imply that the individually detected sources are not extreme cases, Si ii* halos must be common for UV-bright galaxies at –.
@@
The bottom row of Fig. 4 shows the results of stacking all the Si ii* narrowbands and corresponding UV continuum bands for the 14 sources (three lines; in total 22 images). We find that the mean-stacked Si ii* deviates by more than from the mean-stacked UV continuum for six data points. The median-stacked Si ii* is consistent with the mean-stacked Si ii*, and the deviation from the normalized continuum is similarly seen with the median-stacked profiles.
@@
We note that removing MID, which is an X-ray detected AGN with extended Si ii* (see Section 3.2), does not change the results and that the other 13 sources do not have a counterpart in the deep X-ray catalog (Luo et al. 2017). We also note that stacking without individually-detected Si ii* halos results in nondetections of extended Si ii*. As shown in Fig. 2, all the individually-detected sources have a bright (and also a bright ) and are located at relatively low redshifts. It is not surprising that removing such five sources results in nondetections. If the origin of the Si ii* is continuum pumping, as we expect, it then should be more challenging to detect Si ii* halos without these bright sources. We also stack various subsamples including all the sources without individually-detected halos and get nondetections of extended Si ii* emission. Future instruments such as BlueMUSE (Richard et al. 2019) or a larger sample compiling MUSE and KCWI archive data will allow us to investigate the dependence of Si ii* halos on sample properties such as (see Section 4.3). It will enable us to draw a firm conclusion of whether Si ii* halos are indeed common for UV-bright galaxies at – or not.